OpenStudy (anonymous):

The slope of the line passing through the points (2, 7) and (-4, 8) is

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

-1/2 ?

hartnn (hartnn):

what you got for x1 - x2 ?

hartnn (hartnn):

because -1/2 is incorrect.

OpenStudy (anonymous):

let me check again

OpenStudy (anonymous):

Im guessing but would it be -6

OpenStudy (anonymous):

well 2--4 equals 6 then 7-8=1 so x=-6

OpenStudy (anonymous):

i meant 2-(-4)

OpenStudy (anonymous):

ok,thanks:)

OpenStudy (anonymous):

welcome :)

hartnn (hartnn):

yes, so you got -1/6, right ?

OpenStudy (anonymous):

no hun its -6

hartnn (hartnn):

how?

hartnn (hartnn):

i was asking for slope m =-1/6

OpenStudy (anonymous):

please explain:)

OpenStudy (anonymous):

no i is -6/1 which is the same as -6

OpenStudy (anonymous):

because anything over one is itself

hartnn (hartnn):

ok, let me explain, y1-y2 = 7-8 = -1 x1-x2 = 2-(-4) = 6 m = -1/6

OpenStudy (anonymous):

so would it be -1/6/

hartnn (hartnn):

yes

OpenStudy (anonymous):

you are both making it complicated on your selfs. Ill explain

OpenStudy (anonymous):

so what would The slope of the line whose equation is 3y = 2x - 3 be?

hartnn (hartnn):

to get slope of 3y = 2x - 3 divide all terms by 3 to isolate y first

OpenStudy (anonymous):

y2-y1/x2-x1 okay your to points are (2, 7) and (-4, 8) okay well your 2 is your y1, the -4 is your y2 so you take 2 subtract your negative four and anything subtracted by a negative is a positive just look at the drawing i made :) hope it helps